A hot air balloon is filled with 1.37x10^6 of an ideal gas on a cool morning (11C). The air is heated to 121C . What is the volume of the air in the balloon after it is heated? Assume that none of the gas escapes from the balloon.
Answer #1

Answer:

According Charles law, the volume of the gas at constant pressure is directly proportional to absolute temperature.

That means V/T=constant

Or V1/T1=V2/T2

Given V1=1.37 x 10^6 L, T1=11°C=11+273 K=284 K

V2=? T2=121°C=121+273 K=394 K.

Therefore

V2=V1T2/T1=(1.37 x 10^6 L x 284 K)/(394 K)

V2= 1.9 x 10^6 L.

The volume of air after heated=1.9 x 10^6 L.

Please let me know if you have any doubt. Thanks